# Translating predictive distributions into informative priors

### v3

@@ 3.2 Priors from model-derived quantities
assuming that the columns of $\boldsymbol{X}$ have been centred. Our aim is to use our knowledge of $R^{2}$ to set suitable priors for the regression coefficients $\beta$. This idea was the inspiration for a class of shrinkage priors (Zhang & Bondell, 2018; Zhang et al., 2022), but we would like to make this idea applicable to a wider selection of prior structures.
@@ 3.2.1 Evaluation setup and tuning parameters
We fix $n=50$ and $p=80$ with entries in $\boldsymbol{X}$ drawn from a standard Gaussian distribution, and assess replicability using 10 independent runs for each prior and target. The support $\Lambda$ for the hyperparameters is defined in Supplement S5.1. We use $S=10^{4}$ prior predictive samples, $I=5\times 10^{3}$ importance samples from a $\text{Uniform}(0,1)$, and use both $d^{\text{AD}}$ and $d^{\text{CvM}}$ as discrepancy functions. We employ $N_{\text{CRS2}}=1000$ iterations, and subsequently perform both single and multi-objective Bayesian optimisation for $N_{\text{batch}}=1$ batch of $N_{\text{BO}}=150$ iterations, using $N_{\text{design}}=50$ points from the first stage. The single objective approach illustrates that differences in flexibility between priors also induce differences in uniqueness, and highlights issues in choosing a prior for the additive noise parameter $\sigma^{2}$. Choosing $\kappa$ is challenging in the multi-objective approach, as its value should depend on the target, the discrepancy function, and the prior. These dependencies result in 96 possible choices of $\kappa$, which is an infeasible number of choices to make in this example. Instead we fix $\kappa=0.5$ for all multi-objective settings. We use the secondary objective (4), except for quantities where the standard deviation is undefined for some $\lambda\in\Lambda$, for which we use a robust scale estimator (Rousseeuw & Croux, 1993).
@@ 3.2.2 Results
We evaluate faithfulness by inspecting the densities $\text{p}(R^{2}\mid\lambda^{*})$ and $\text{t}(R^{2})$ for the various targets (all distributions in this example have corresponding densities). A selected subset of the pairs of $(s_{1},s_{2})$ values are displayed in Figure 7 (complete results are in Supplement S5.2). The Gaussian prior is universally poorly faithful. Both shrinkage priors perform better in cases where one of $s_{1}$ or $s_{2}$ is less than 1, with the regularised horseshoe performing better for the $s_{1}=s_{2}>1$ cases. Interestingly, the results are not symmetric in $s_{1}$ and $s_{2}$; the Dirichlet-Laplace prior is able to match the $s_{1}=3,s_{2}=0.69$ target well, with many of regularised horseshoe replicates performing poorly; whilst the relative performance is reversed for $s_{1}=0.69,s_{2}=3$ (see Figure 17). There is also perceptibly more variability in the regularised horseshoe replicates, which suggests the optimisation problem is more challenging and the predictive discrepancy objective is noisier. The multi-objective generally produces more variable sets of optima, which is expected, as it is a more difficult optimisation problem but we do not allow it additional computational resources. There is little visible difference between the two discrepancy functions. Finally, as the values of $s_{1}$ and $s_{2}$ increase, the faithfulness of the shrinkage priors generally decreases. Across the full set of simulations, the regularised horseshoe is evidently the most flexible.
@@
The hyperparameters $(a_{1},b_{1})$ for the additive noise variance $\sigma^{2}$ are highly variable across replications for almost all prior/target combinations. This reflects the anticipated lack of uniqueness when incorporating such hyperparameters. It is particularly striking for the Dirichlet-Laplace prior when $s_{2}\in\{0.33,0.69\}$, where we observe consistent and excellent fits/faithfulness to the target, but these do not correspond to replicable estimates for $(a_{1},b_{1})$. These settings are also interesting as the choice of single or multi-objective approach greatly impacts the optimum values of $a_{1}$ and $b_{1}$. Faithfulness of the multi-objective optima, illustrated in Figure 7, are not appreciably worse than the single objective approach, but the inclusion of $\sigma^{2}$ into the secondary objective has resulted in optimal values of $a_{1}$ and $b_{1}$ that maximise the dispersion of the marginal prior (i.e. small $a_{1}$ and large $b_{1}$). Asymptotic results are also known for the Gaussian prior, and in Supplement S5.3 we further assess replicability by benchmarking our optimisation process against suitable ‘true’ (asymptotically) values.
Our optimisation procedure has minimised $\log(D(\lambda))$ using both the AD and CvM discrepancy functions. The former places extra emphasis on matching the tails of the target, and thus the Regularised Horseshoe values in the top row of Figure 6 differ from our expectations given the results in the top two rows of Figure 7. Take, for example, the $s_{1}=3,s_{2}=0.69$ case. It is plainly evident from Figure 7 that the regularised horseshoe prior provides a better fit to the target distribution at $\boldsymbol{\lambda}^{*}_{\text{HS}}$, and yet the corresponding $\log(D(\boldsymbol{\lambda}_{\text{HS}}^{*}))$ values in the top row of Figure 6 suggest that it is considerably worse that the Gaussian prior at $\boldsymbol{\lambda}_{\text{GA}}^{*}$. To reconcile this apparent contradiction, we inspect $\log(D(\lambda))$ at the optima computed using the Cramér-Von Mises discrepancy function. These values are displayed in the bottom row of Figure 6, whose values closely match our expectations given Figure 7. Given the range of behaviours of $\text{p}(R^{2}\mid\lambda^{*})$ for all the optima, we can conclude that AD more heavily penalises over-estimation of the tails of $\text{p}(R^{2}\mid\lambda^{*})$ than under-estimation. This does not discount it as an optimisation objective, but does complicate comparisons between competing priors.
@@ 3.3 A human-aware prior for a human growth model
with $\varepsilon_{m}\sim\text{N}(0,\sigma^{2}_{y})$. Some constraints are required to identify this model and ensure its physical plausibility: specifically, we require $0<h_{0}<h_{1}$ and $0<s_{0}<s_{1}$. To satisfy these constraints, we parameterise in terms of $\delta_{h}=h_{1}-h_{0}$ and $\delta_{s}=s_{1}-s_{0}$, which results in $(h_{0},\delta_{h},s_{0},\delta_{s})$ all sharing the same positivity constraint. We also constrain $\gamma$ such that $\gamma\in(\min_{m}(t_{m}),\max_{m}(t_{m}))$. Even with these constraints the the denominator of the fraction can be very small, yielding negative heights, meaning the model is not plausible for all parameter values. Furthermore, the model is poorly behaved under a flat prior, so prior information is required to stabilise and/or regularise the posterior.
We thus seek in this example to specify priors congruent with two specific target prior predictive distributions. We choose $\text{LogNormal}(\mu_{q},s^{2}_{q})$ priors for each of the $q=1,\ldots,5$ elements of $\theta=(h_{0},\delta_{h},s_{0},\delta_{s},\gamma)$, and seek to identify the optimal values of $\lambda=\left(\mu_{q},s^{2}_{q}\right)_{q=1}^{5}$ (see Supplement S6.1 for $\Lambda$). We fix the prior $\sigma_{y}\sim\text{LogNormal}(0,0.2^{2})$ to avoid identifiability problems (Section 3.2). We suppose our data originate from humans, uniformly distributed between ages 2 and 18, and evenly split between sexes. We first consider a covariate-independent prior predictive density $\text{t}(Y)$ with corresponding CDF $\text{T}(Y)$ for human heights across the entire age-range, derived by summarising external data. This target (Figure 10) is a mixture of 3 gamma densities specified to approximate the external data, which is multimodal due to the fact that humans grow in spurts. We also consider a covariate-specific $\text{T}(Y\mid X_{r})$, specifying Gaussian height distributions at ages $X_{r}\in(2,8,13,18)$ (see Figure 11, and Supplement S6.2 for details).
@@
Hartmann et al. (2020) also considered this example, but elicited 6 predictive quantiles at ages $t=(0,2.5,10,17.5)$, as opposed to entire predictive distributions at ages $t=(2,8,13,18)$ as in our covariate-specific approach. We use different ages because the model is stated to be accurate for ages $\geq 2$ (Preece & Baines, 1978). Hartmann et al. (2020) also include a noise parameter in their definition of $\theta$. The exact interpretation of this parameter is complicated by their choice of Weibull likelihood, rendering the distribution of the measurement errors sensitive to conditional mean of the model (this is still the case despite their choice of Weibull parameterisation). Finally, Hartmann et al. (2020) elicit quantiles from 5 different users and report an estimated $\lambda^{*}$ for each user. These estimates (reproduced in Supplement S6.3) allow us to compare optimal the selected priors $\text{p}(\theta\mid\lambda^{*})$
For both targets we obtain $\lambda^{*}$ using both the single objective and multi-objective optimisation processes. We use only the Cramér-Von Mises discrepancy in this example, because we were not able to reliably compute the AD discrepancy due to numerical instabilities. We use $S=5\times 10^{4}$ samples from $\text{p}(Y\mid\lambda)$ and likewise $S_{r}=5\times 10^{4}$ samples from $\text{p}(Y\mid\lambda,X_{r})$ for each of the 4 values of $X_{r}$. We use $I=5\times 10^{3}$ and $I_{r}=5\times 10^{3}$ importance samples, with $N_{\text{CRS2}}=2000$ CRS2 iterations, $N_{\text{batch}}=5$ Bayesian optimisation batches each of $N_{\text{BO}}=250$ iterations, and carry forward $N_{\text{design}}=50$ points per batch. We assess replicability using 30 independent runs of each objective/target pair.
@@ 3.3.2 Results
We consider $\kappa\in\mathcal{K}=\{0.05,0.1,\ldots,0.5\}$ for the covariate-specific targets. There is notable inter-replicate variability in the Pareto frontiers for the selection of values $\kappa\in\{0.1,0.2,0.3\}$ shown in Figure 9, due to the stochastic properties of the global optimisers we employ, with some replicates totally dominated by other replicates. Following our ‘minimum variability across replicates’ heuristic (Section 2.4), we select $\kappa=0.2$ for the covariate-specific target and $\kappa=0.15$ for the covariate-independent target (See Supplement S6.4). With these values the optimisation procedure identifies optimal $\lambda^{*}$ values with reasonably, but not entirely, consistent predictive discrepancies across replicates (see Supplement S6.5).
@@
Figure 10: The covariate-independent marginal target density $\text{t}(Y)$ (red) and prior predictive densities $\text{p}(Y\mid\lambda^{*})$ for each of the 30 replicates (blue lines).
@@
Figure 10 displays the target and prior predictive density estimates in the covariate-independent case. The multi-objective replicates are obtained after $\kappa=0.2$ is chosen. We see that introducing the secondary objective produces estimates of $\lambda^{*}$ that are congruent with estimates from the single objective case, though the estimates are more variable. Both single and multi-objective approaches result in reasonably, but not entirely, faithful densities for $\text{p}(Y\mid\lambda^{*})$. However, most optimum priors seem to accumulates additional probability surrounding $Y=h_{1}\approx 155$, resulting in individual trajectories attaining their adult height $h_{1}$ for younger than expected ages $t$ (which we will later confirm in Figure 12).
Figure 11: Covariate-specific target densities $\text{t}(Y\mid X_{r})$ (red lines) and prior predictive densities $\text{p}(Y\mid\lambda^{*},X_{r})$ for each of the 30 replicates (blue lines).
For the covariate-specific target (Figure 11), the secondary objective introduces in some replicates outlying estimates for $\text{p}(Y\mid\lambda^{*},X_{r})$, most clearly visible for $X_{1}=2$ and $X_{2}=8$. Both the single and multi-objective approaches struggle to match the prior predictive distribution at all ages, with consistently poorer faithfulness for $X_{1}=2$. Empirically, it does not seem possible to match all four supplied target prior predictive distributions simultaneously, given the mathematical structure of the model. Lastly, because $\text{t}(Y\mid X_{1}=2)$ is substantially narrower than the other targets, it is optimal, under the Cramér-Von Mises discrepancy, to select wider priors better matching the older age target distributions.
@@
Figure 12: Mean (solid lines) and 95% intervals (grey regions) for prior predictives $\text{p}(h(t;\theta)\mid\lambda^{*})$ for covariate-independent and covariate-specific targets for the multi- and single-objective settings; and for each replicate from Hartmann et al. (with 75% intervals). The y-axis is truncated to $(50,200)$. The red lines are 95% intervals for $\text{t}(Y\mid X_{r})$.
@@
Both the covariate-independent and covariate-specific targets yield plausible mean growth trajectories (Figure 12). However, covariate-independent priors are highly uncertain, resulting in implausible heights having a priori support. It is interesting to note that the covariate-specific target has similar levels of uncertainty across all ages, further suggesting that the model is too inflexible to simultaneously match all the targets when they have varying variance, as here. All 5 of the priors from Hartmann et al. (2020), for a narrower uncertainty interval, are implausible in both shape and width when viewed on this scale. It also seems unlikely that these priors accurately reflect the information provided by the experts in Hartmann et al. (2020), but this information is not reported.
@@
Figure 13: Presence/absence of Stan warnings for all individuals (columns) in the FDA package growth data and replicate prior estimates (rows). Each replicate corresponds to a run of the optimisation process and thus a different prior (except for Flat).
@@
The different priors produce a widely ranging proportion of warning messages in Stan (Figure 13). The flat prior produces the most warnings, with some individuals particularly prone to warning messages, suggesting that their data are relatively uninformative. The Hartmann priors produce a moderate number of warnings, with some priors less prone to produce warnings (replications 1 and 5) than others for this dataset. The covariate-specific approach produces fewer warnings than the covariate-independent approach in both the single- or multi-objective cases. This reflects the additional information available in the covariate-specific setting, and that this information results in more plausible priors. Some specific replications of the covariate-independent approach produce many warnings, suggesting these priors are inappropriate for many individuals.
@@
Figure 14: Priors (blue) for $h_{0}$ with covariate-independent and covariate-specific targets; a flat prior scenario (prior not shown); and Hartmann et al. (2020). The corresponding posteriors for each prior for individual $n=26$ is shown in red.
@@
The priors for $h_{0}$ exhibit substantial variability across replicates (Figure 14; see Supplement S6.6 for all $\theta$). There appear to be two distinct unimodal priors for $h_{0}$ with similar loss, suggesting that $\text{T}(Y\mid\boldsymbol{X})$ does not provided enough information to uniquely determine a prior distribution. However both priors are significantly broader than the Hartmann et. al. priors. Figure 14 also shows the posteriors for these parameters when using the (uninformative and thus challenging) data from individual $n=26$. The flat prior produces a multimodal posterior for our parameters of interest, demonstrating the lack of stability when computing the posterior and the need for prior information. As a result, the posteriors are under all methods strongly depend on the prior distribution used, which as noted is not stable under any method here. However both of our priors produce posteriors for $\delta_{s}$ with no significant mass above 2 (Supplement S6.6), as is desirable, because this corresponds physiologically implausible growth spurts that are unsupported by the data.
@@
In summary, the priors estimated by our procedure in this example are broadly faithful to the supplied information, except in the covariate-specific case where model inflexibility prevents matching both $t=2$ and $t=18$ targets simultaneously. The covariate-specific, multi-objective method appears the most useful prior, but is arguably over concentrated, which occasionally prevents the model from fitting the data well, although all our priors successfully regularise the posterior sufficiently to enable accurate posterior sampling. Our approach does not produce a unique prior, although the secondary objective leads to a small improvement in uniqueness (see Supplement S6.6). However, some of this non-uniqueness may be attributable to imperfect replicability of the optimisation.

### v4

@@ 3.2 Priors from model-derived quantities
assuming that the columns of $\boldsymbol{X}$ have been centred. Our aim is to use our knowledge of $R^{2}$ to set suitable priors for the regression coefficients $\beta$ conditional on $\boldsymbol{X}$. This idea was the inspiration for a class of shrinkage priors (Zhang & Bondell, 2018; Zhang et al., 2022), but we would like to make this idea applicable to a wider selection of prior structures.
@@ 3.2.1 Evaluation setup and tuning parameters
We fix $n=50$ and $p=80$ with entries in $\boldsymbol{X}$ drawn from a standard Gaussian distribution, and assess replicability using 10 independent runs for each prior and target. The support $\Lambda$ for the hyperparameters is defined in Supplement S6.1. We use $S=10^{4}$ prior predictive samples, $I=5\times 10^{3}$ importance samples from a $\text{Uniform}(0,1)$, and use both $d^{\text{AD}}$ and $d^{\text{CvM}}$ as discrepancy functions. We employ $N_{\text{CRS2}}=1000$ iterations, and subsequently perform both single and multi-objective Bayesian optimisation for $N_{\text{batch}}=1$ batch of $N_{\text{BO}}=150$ iterations, using $N_{\text{design}}=50$ points from the first stage. The single objective approach illustrates that differences in flexibility between priors also induce differences in uniqueness, and highlights issues in choosing a prior for the additive noise parameter $\sigma^{2}$. Choosing $\kappa$ is challenging in the multi-objective approach, as its value should depend on the target, the discrepancy function, and the prior. These dependencies result in 96 possible choices of $\kappa$, which is an infeasible number of choices to make in this example. Instead we fix $\kappa=0.5$ for all multi-objective settings. We use the secondary objective (5), except for quantities where the standard deviation is undefined for some $\lambda\in\Lambda$, for which we use a robust scale estimator (Rousseeuw & Croux, 1993).
@@ 3.2.2 Results
We evaluate faithfulness by inspecting the densities $\text{p}(R^{2}\mid\lambda^{*})$ and $\text{t}(R^{2})$ for the various targets (all distributions in this example have corresponding densities). A selected subset of the pairs of $(s_{1},s_{2})$ values are displayed in Figure 7 (complete results are in Supplement S6.2). The Gaussian prior is universally poorly faithful. Both shrinkage priors perform better in cases where one of $s_{1}$ or $s_{2}$ is less than 1, with the regularised horseshoe performing better for the $s_{1}=s_{2}>1$ cases. Interestingly, the results are not symmetric in $s_{1}$ and $s_{2}$; the Dirichlet-Laplace prior is able to match the $s_{1}=3,s_{2}=0.69$ target well, with many of regularised horseshoe replicates performing poorly; whilst the relative performance is reversed for $s_{1}=0.69,s_{2}=3$ (see Supplement Figure 16). There is also perceptibly more variability in the regularised horseshoe replicates, which suggests the optimisation problem is more challenging and the predictive discrepancy objective is noisier. The multi-objective approach generally produces more variable sets of optima, which is expected, as it is a more difficult optimisation problem but we do not allow it additional computational resources. There is little visible difference between the CvM and AD discrepancy functions. Finally, as the values of $s_{1}$ and $s_{2}$ increase, the faithfulness of the shrinkage priors generally decreases. Across the full set of simulations, the regularised horseshoe is evidently the most flexible.
@@
The hyperparameters $(a_{1},b_{1})$ for the additive noise variance $\sigma^{2}$ are highly variable across replications for almost all prior/target combinations. This reflects the anticipated lack of uniqueness when incorporating such hyperparameters. It is particularly striking for the Dirichlet-Laplace prior when $s_{2}\in\{0.33,0.69\}$, where we consistently attain faithfulness but no replicability in estimates for $(a_{1},b_{1})$. These settings are also interesting as the choice of single or multi-objective approach greatly impacts the optimum values of $a_{1}$ and $b_{1}$. Faithfulness of the multi-objective optima, illustrated in Figure 7, are not appreciably worse than the single objective approach, but the inclusion of $\sigma^{2}$ into the secondary objective has resulted in optimal values of $a_{1}$ and $b_{1}$ that maximise the dispersion of the marginal prior (i.e. small $a_{1}$ and large $b_{1}$). Asymptotic results are also known for the Gaussian prior, and in Supplement S6.3 we further assess replicability by benchmarking against (asymptotically) ‘true’ values.
Our optimisation procedure has minimised $\log(D(\lambda))$ using both the AD and CvM discrepancy functions. The former places extra emphasis on matching the tails of the target, and thus the Regularised Horseshoe values in the top row of Figure 6 differ from our expectations given the results in the top two rows of Figure 7. Take, for example, the $s_{1}=3,s_{2}=0.69$ case. It is plainly evident from Figure 7 that the regularised horseshoe prior provides a better fit to the target distribution at $\boldsymbol{\lambda}^{*}_{\text{HS}}$, and yet the corresponding $\log(D(\boldsymbol{\lambda}_{\text{HS}}^{*}))$ values in the top row of Figure 6 suggest that it is considerably worse that the Gaussian prior at $\boldsymbol{\lambda}_{\text{GA}}^{*}$. To reconcile this apparent contradiction, we inspect $\log(D(\lambda))$ at the optima computed using the CvM discrepancy function. These values are displayed in the bottom row of Figure 6, whose values closely match our expectations given Figure 7. Given the range of behaviours of $\text{p}(R^{2}\mid\lambda^{*})$ for all the optima, we can conclude that AD more heavily penalises over-estimation of the tails of $\text{p}(R^{2}\mid\lambda^{*})$ than under-estimation. This does not discount it as an optimisation objective, but does complicate comparisons between competing priors.
@@ 3.3 A human-aware prior for a human growth model
with $\varepsilon_{m}\sim\text{N}(0,\sigma^{2}_{y})$. Some constraints are required to identify this model and ensure its physical plausibility: specifically, we require $0<h_{0}<h_{1}$ and $0<s_{0}<s_{1}$. To satisfy these constraints, we parameterise in terms of $\delta_{h}=h_{1}-h_{0}$ and $\delta_{s}=s_{1}-s_{0}$, which results in $(h_{0},\delta_{h},s_{0},\delta_{s})$ all sharing the same positivity constraint. We also constrain $\gamma$ such that $\gamma\in(\min_{m}(t_{m}),\max_{m}(t_{m}))$. Even with these constraints the denominator of the fraction can be very small, yielding negative heights, meaning the model is not plausible for all parameter values. Furthermore, the model is poorly behaved under a flat prior, so prior information is required to stabilise and/or regularise the posterior.
We thus seek in this example to specify priors congruent with two specific target prior predictive distributions. We choose $\text{LogNormal}(\mu_{q},s^{2}_{q})$ priors for each of the $q=1,\ldots,5$ elements of $\theta=(h_{0},\delta_{h},s_{0},\delta_{s},\gamma)$, and seek optimal values of $\lambda=\left(\mu_{q},s^{2}_{q}\right)_{q=1}^{5}$ (see Supplement S7.1 for $\Lambda$). We fix the prior $\sigma_{y}\sim\text{LogNormal}(0,0.2^{2})$ to avoid uniqueness problems (Section 3.2). We suppose both sex and age (between ages 2 and 18) are uniformly distributed in our data. We first consider a covariate-independent prior predictive density $\text{t}(Y)$ with corresponding CDF $\text{T}(Y)$ for height across the entire age-range, derived by summarising external data. This target (Figure 10) is a mixture of 3 gamma densities specified to approximate the external data, which is multimodal due to the fact that humans grow in spurts. We also consider a covariate-specific $\text{T}(Y\mid X_{r})$, specifying Gaussian height distributions at ages $X_{r}\in(2,8,13,18)$ (see Figure 22, and Supplement S7.2 for details).
@@
Hartmann et al. (2020) also considered this example, but elicited 6 predictive quantiles at ages $t=(0,2.5,10,17.5)$, as opposed to entire predictive distributions at ages $t=(2,8,13,18)$ as in our covariate-specific approach. We use different ages because the model is stated to be accurate for ages $\geq 2$ (Preece & Baines, 1978). Hartmann et al. (2020) also include in their definition of $\theta$ a noise parameter; the distribution of this depends on the conditional mean of the model due to the Weibull likelihood adopted by Hartmann et al. (2020). Finally, Hartmann et al. (2020) elicit quantiles from 5 different users and report an estimated $\lambda^{*}$ for each user. These estimates (reproduced in Supplement S7.3) allow us to compare optimal the selected priors $\text{p}(\theta\mid\lambda^{*})$.
We obtain $\lambda^{*}$ for both targets using both single- and multi-objective optimisation processes. We use the CvM discrepancy, and both forward and reverse KL discrepancies (numerical instability prevented use of the AD discrepancy). We use $S=5\times 10^{4}$ samples from $\text{p}(Y\mid\lambda)$ and likewise $S_{r}=5\times 10^{4}$ samples from $\text{p}(Y\mid\lambda,X_{r})$ for each of the 4 values of $X_{r}$. We use $I=5\times 10^{3}$ and $I_{r}=5\times 10^{3}$ importance samples for the CvM discrepancy, and the same number of samples for estimating the relevant Gaussian parameters in the KL approximation. Lastly, all settings use $N_{\text{CRS2}}=2000$ CRS2 iterations, $N_{\text{batch}}=5$ Bayesian optimisation batches each of $N_{\text{BO}}=250$ iterations, and carry forward $N_{\text{design}}=50$ points per batch. We assess replicability using 30 independent runs of each objective/target pair.
@@ 3.3.2 Results
We consider target- and discrepancy-specific ranges $\kappa\in\mathcal{K}$ for the multi-objective settings, and follow our ‘minimum variability across replicates’ heuristic (Section 2.4) to select optimum $\kappa^{*}$ values (listed in Supplement S7.4). There is notable inter-replicate variability in the Pareto frontiers at the optimal values $\kappa^{*}$ (Figure 9), due to the stochasticity of our two-stage optimisation approach, with some replicates totally dominated by other replicates. The predictive discrepancies for the corresponding optimal $\lambda^{*}$ values are reasonably, but not entirely, consistent across replicates (see Supplement S7.6).
@@
Figure 10: The covariate-independent marginal target density $\text{t}(Y)$ (red) and prior predictive densities $\text{p}(Y\mid\lambda^{*})$ for each of the 30 replicates (blue).
@@
Figure 10 displays the target and prior predictive density estimates in the covariate-independent case. The multi-objective replicates are obtained after $\kappa^{*}$ is chosen. We see that introducing the secondary objective produces estimates of $\lambda^{*}$ that are congruent with the single objective case, but are more variable. Both single and multi-objective approaches result in reasonably, but not entirely, faithful densities for $\text{p}(Y\mid\lambda^{*})$, though the KL-based discrepancies are notably less faithful. However, most optimum priors seem to accumulate additional probability surrounding $Y=h_{1}\approx 155$ (for the CvM discrepancy) or $\approx 125$ (for the KL discrepancies), resulting in individual trajectories attaining their adult height $h_{1}$ for younger than expected ages $t$ (which we will later assess in Figure 11). We similarly assess faithfulness to the covariate-specific target in Supplement S7.7, noting that the reverse-KL exhibits over-concentration compared to the other discrepancies (which is to be expected, see Minka (2005)).
@@
Figure 11: Mean (solid lines) and 95% intervals (grey regions) for the prior mean $\text{p}(h(t;\theta)\mid\lambda^{*})$, for covariate-independent and covariate-specific targets in the multi- and single-objective settings, for all discrepancies; and for the Hartmann priors (with 75% intervals). The y-axis is truncated to $(50,200)$. The red lines are 95% intervals for $\text{t}(Y\mid X_{r})$.
@@
Figure 11 shows that both the covariate-independent and covariate-specific targets yield plausible mean growth trajectories for the CvM discrepancy, however only the covariate-specific target does so for the KL-based discrepancies. The covariate-independent priors are more uncertain, resulting in implausible heights having a priori support. The covariate-specific priors have similar levels of uncertainty across all ages, further suggesting that the model is too inflexible to simultaneously match all the covariate-specific targets, which have varying variance. All 5 of the priors from Hartmann et al. (2020), for a narrower uncertainty interval, are implausible in both shape and width when viewed on this scale. It also seems unlikely that these priors accurately reflect the information provided by the experts in Hartmann et al. (2020), but this information is not reported.
@@
Figure 12: Presence/absence of Stan warnings for all individuals (columns) in the growth data and replicate prior estimates (rows). Each replicate corresponds to a run of the optimisation process (CvM discrepancy) and thus a different prior (except for flat).
@@
The different priors produce a widely ranging proportion of warning messages in Stan (Figure 12). The flat prior produces the most warnings, with some individuals particularly prone to warning messages, suggesting that their data are relatively uninformative. The Hartmann priors produce a moderate number of warnings, with some priors less prone to produce warnings (replications 1 and 5) than others for this dataset. Using the CvM discrepancy, the covariate-specific approach produces fewer warnings than the covariate-independent approach in both the single or multi-objective cases. This reflects the additional information available in the covariate-specific setting that results in more informative and plausible priors. Some specific replications of the covariate-independent approach produce many warnings, suggesting these priors are inappropriate for many individuals.
@@
Figure 13: Priors (blue) for $h_{0}$ and corresponding posteriors (red) for individual $n=26$ using either covariate-independent and covariate-specific targets, for the CvM discrepancy; a flat prior scenario (prior not shown); and Hartmann et al. (2020).
@@
The priors for $h_{0}$ exhibit substantial variability across replicates (Figure 13; see Supplement S7.8 for a comparison to the KL discrepancy, and Supplement S7.9 for all $\theta$). Under both covariate-independent and covariate-specific approaches, there are two distinct unimodal priors for $h_{0}$ with similar loss, suggesting that $\text{T}(Y\mid\boldsymbol{X})$ does not provided enough information to uniquely determine a prior distribution. However both priors are significantly broader than the Hartmann et. al. priors. Figure 13 also shows the posteriors for these parameters when using the (uninformative and thus challenging) data from individual $n=26$. The posterior sampler, in the flat prior setting, unreliably locates and adapts to the posterior, resulting in the varied posterior estimates visible in Figure 13. Conversely, any single replicate from the informative prior approaches provides enough regularization to ensure consistent posterior sampling for that particular replicate. Given this sensitivity to prior information, the posteriors strongly depend on the prior distribution used, which as noted is not stable under any method here. However, our priors produce posteriors for $\delta_{s}$ with almost all mass below 2 (Supplement S7.9); this is desirable, because $\delta_{s}>2$ corresponds to physiologically implausible growth spurts that are unsupported by the data.
@@
In summary, the priors estimated by our procedure in this example are broadly faithful to the supplied information, except in the covariate-specific case where model inflexibility prevents matching both $t=2$ and $t=18$ targets simultaneously, and in the covariate-independent case when either KL discrepancy is used. The covariate-specific, multi-objective method appears the most useful prior, but is arguably over concentrated, which occasionally prevents the model from fitting the data well, although all our priors successfully regularise the posterior sufficiently to enable accurate posterior sampling. Our approach does not produce a unique prior, although the secondary objective leads to a small improvement in uniqueness (see Supplement S7.9). However, some of this non-uniqueness may be attributable to imperfect replicability of the optimisation.